40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 9 10:57:06 2021, Paco, Anchal, Summary, IMC, XARM lock and POX spectra 20210309_POX11_Spec_XARMLocked.pdf20210309_XARM_Locked.tar.gz
    Reply  Tue Mar 9 14:37:26 2021, gautam, Summary, SUS, PRM suspension PRMtrip.png
       Reply  Tue Mar 9 15:22:56 2021, Koji, Summary, SUS, PRM suspension 
Message ID: 15884     Entry time: Tue Mar 9 10:57:06 2021     Reply to this: 15887
Author: Paco, Anchal 
Type: Summary 
Category: IMC 
Subject: XARM lock and POX spectra 

[Paco, Anchal]

- Upon arrival, MC is locked, and we can see light in MON5 (PRM) (usually dark).

# XARM locking
- Read through "XARM POX" script (path='/cvs/cds/rtcds/caltech/c1/burt/c1configure/c1configureXarm')
- Before running the script, we noticed the PRM watchdog is down, so we manually repeat the procedure from last time, but see more swinging even though the watchdog is active.
- Run a reEnablePRMWatchdogs.py script (a copy of reEnableWatchdogs.py with optics=['PRM']), which had the same effect. 
- We manually disable the watchdog to recover the state we first encountered, and wait for the beam in MON5 to come to rest.
    - The question is; is it fine to lock Xarm with PRM watchdog down?
    - To investigate this, we look at the effect of the offset on the unwatchdog-PRM.
    - Manually change 'PRM_POS_OFFSET' to 200, and -800 (which is the value used in the script) with no effect on the PRM swinging.
- Moving on, run IFO > CONFIGURE > ! (X Arm) > RESTORE XARM (XARM POX), and ... success.

# MC-POX noise spectra
- With XARM locked, open diaggui and take spectra for C1:LSC-POX11_I_ERR_DQ, C1:LSC-POX11_Q_ERR_DQ, C1:IOO-MC_F_DQ
- Lost XARM lock while we were figuring out unit conversions...
    - Assuming 2.631e-13 m/counts (6941) and using 37.79 m (arm length), 1064.1 nm wavelength, we get a calibration factor of 2.631e-13 * c / (2*L*lambda) ~ 0.9809 Hz/count 
    - (FAQ?, how to find/compute/measure the correct calibration factors?)
- Relock XARM, retake spectra. Attachment 1 has plots for POX11_I/Q_ERR_DQ spectrum (cts/rtHz, we couldn't find relevant calibration) and MC_F_DQ in (Hz/rtHz from referring to 15576, we couldn't get the units to show on y scale.)

# MC-POY noise spectra (attempt)
- Now, run IFO > CONFIGURE > ! (Y Arm) > RESTORE YARM (YARM POY), and XARM locks (why?)
    - Could PRM watchdog being down be the cause? 
- Try C1ASS > (YARM) ! More Scripts > ON, and looked at YARM PIT/YAW striptool. 
- C1ASS > (YARM) ! Freeze Outputs, then OFF
- Go back to IFO > CONFIGURE > ! (Y Arm) > Align YARM  (ASS ON: Unfreeze), try running this then Freeze, then OFF Zero Outputs.
- Try RESTORE YARM (POY) again, still not working.
- Try RESTORE YARM ALS, then try again after opening the shutter, but also fail to lock AUX.
    - Is the PRM WD behind some evil misalignment? Will move forward with XARM bc it is happy.

# ARM locking
- Attempted the IFO > CONFIGURE > ! (X Arm) > RESTORE Xarm (XARM ALS) but green failed to lock and we lost XARM lock.
- Try to recover XARM lock... success. It's nice to have a (repeatable) checkpoint.
- Attempt YARM lock. Not successful. It just seems like the lock Triggers are not raised (misalignment?)
    - From C1SUS_ETMY, try changing the bias "C1:SUS-ETMY_YAW_OFFSET" manually to reduce the OPLEV_YERROR. Changed from -47 to -57.
    - Retry YARM lock script... no luck
    - From C1SUS_PRM, try changing the bias "C1:SUS-PRM_PIT_OFFSET" manually to reduce OPLEV errors. Changed from 34 to 22 with no effect, then realized the coil outputs are disabled because the WD is down...
    - So we do the following BIAS changes "C1:SUS-PRM_PIT_OFFSET" = 34 > 770 and "C1:SUS-PRM_YAW_OFFSET" = 134 > -6
    - Enable all Coil Outputs, turn WD to Normal, turn OPLEVs ON, (this time the beam does not swing like crazy).
    - Fine tune BIASes "C1:SUS-PRM_PIT_OFFSET" = 770 > 805  and "C1:SUS-PRM_YAW_OFFSET" = -6 > 65
        - Saw YARM locking briefly, then unlocking, but we stopped once the OPLEV_ERRs no longer overloaded (from magnitudes > 50 to ~ 40).
- Retry YARM lock... no luck
    - From C1SUS_ETMY, try changing the bias "C1:SUS-ETMY_PIT_OFFSET" from -1 to 6. 

Stop for the day. Leave XARM locked, MC locked. 

Attachment 1: 20210309_POX11_Spec_XARMLocked.pdf  53 kB  | Hide | Hide all
20210309_POX11_Spec_XARMLocked.pdf
Attachment 2: 20210309_XARM_Locked.tar.gz  1.695 MB
ELOG V3.1.3-